OpenStudy (anonymous):

Last month you had $698 in your savings account and received $5.58 in interest. This month you have been able to save a total of $1406. Assuming the rate has stayed the same, how much interest will you earn? A. $11.24 B. $2.01 C. $14.68 D. $8.32

OpenStudy (campbell_st):

I'd use \[5.58 \div 698 = answer\] then multiply the answer by $1406

OpenStudy (texaschic101):

or you can set it up as a proportion.. 698/5.58 = 1406/x -- 698 to 5.58 interest = 1406 to x interest cross multiply (698)(x) = (5.58)(1406) 698x = 7845.48 x = 7845.48/698 x = 11.239 which rounds to $ 11.24 (A)